OpenStudy (anonymous):

Medal The daily adult dosage for a particular drug is 10 mg/kg. How many grams of the drug should be administered daily to a patient weighing 176 lbs? ? g

OpenStudy (danjs):

\[\frac{ 176 lbs }{ 1 }\frac{ 0.454 kg }{ 1 lb }\frac{ 10 mg }{ 1 kg }\]

OpenStudy (anonymous):

change 176lb to kg then multiple 10mg by the kgs to get the amount of mg

OpenStudy (fibonaccichick666):

I prefer using 2.2 lbs to kilogram, but either works

OpenStudy (danjs):

(176*0.454*10) mg

OpenStudy (danjs):

the important thing in these questions i am assuming is to keep track of your units, dimensional analysis

OpenStudy (anonymous):

799.04

OpenStudy (danjs):

that is mg

OpenStudy (danjs):

convert that into grams by \[\frac{ 799.04 g }{ 1 }\frac{ 1 kg }{ 10^3 g}\]

OpenStudy (anonymous):

omg

OpenStudy (luckycoins888):

thats hard

OpenStudy (danjs):

799.04 / 1000 GRAMS

OpenStudy (anonymous):

0.79904

OpenStudy (fibonaccichick666):

@harz360 , do you understand how he developed the equations to modify the lbs to kg? then mg to kg?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

so the answer is 0.79904

OpenStudy (fibonaccichick666):

cool just making sure :)

OpenStudy (anonymous):

is the answer 0.79904